OpenStudy (anonymous):

Which of the following are geometric sequences? Check all that apply. A. 10, 5, 2.5, 1.25, 0.625, 0.3125 B. 5, 10, 20, 40, 80, 160 C. 3, 6, 9, 12 ,15, 18 D. 1, 3, 9, 27, 81

OpenStudy (lukecrayonz):

Arithmetic sequence = adding or subtracting Geometric sequence = division or multiplication

OpenStudy (lukecrayonz):

A. The common "term" is (its actually a ratio) multiplying by 0.5, or dividng by 2 B. 5*2=10*2=20*2=40 So common ratio is 2 C. 3+3=6+3=9+3=12+3=15 So adding 3 is the common term D. 1*3=3*3=9*3=27 so common ratio is multiplying by 3

OpenStudy (anonymous):

so choices A & D would both apply to being geometric seqences?

OpenStudy (ranga):

A, D and one more

OpenStudy (lukecrayonz):

Look at B again, it's common ratio is multiplying by 2.

OpenStudy (anonymous):

oh okay so since it has multiplication A, B, and D would be the answers?

OpenStudy (ranga):

yes.

OpenStudy (ranga):

To find if it is a geometric sequence, just divide 2nd term by first term, 3rd term by 2nd term, etc. and see if the ratio is the same. If it is then you have a geometric series. To figure if a sequence is arithmetic, take the difference of successive terms and see if the differences are the same. If it is then you have an arithmetic sequence.

OpenStudy (ranga):

series should be sequence.

OpenStudy (anonymous):

@ranga so A,B,D?

OpenStudy (ranga):

Yes, A,B & D are geometric sequences. C is an arithmetic sequence.

OpenStudy (anonymous):

thank you :)

OpenStudy (ranga):

you are welcome.